OpenStudy (anonymous):

how many solutions are there to the equation 14x + 2 = 12x

OpenStudy (australopithecus):

I explained how to solve these problems already

OpenStudy (anonymous):

do you know how to solve equations?

OpenStudy (australopithecus):

what are you not getting about it

OpenStudy (australopithecus):

if you dont want to learn how to solve problems such as these just type the problem into https://www.wolframalpha.com/i

OpenStudy (australopithecus):

sorry I mean https://www.wolframalpha.com/

OpenStudy (australopithecus):

it will give you your answer

OpenStudy (anonymous):

still trying to grasp it its been a long time since i took algebra and now i am attending school on-line and it is very challenging!!!

OpenStudy (australopithecus):

I'm not trying to offend, can you at least make an attempt at the problem?

OpenStudy (anonymous):

i am trring a few practice questions

OpenStudy (australopithecus):

Well the first thing you need to do with these question is isolate the x on one side of the equation and the numbers on the otherside

OpenStudy (australopithecus):

This is an equality, so what we do to one side we must do to the other or we will lose the equality. 14x + 2 = 12x Subtract 14x from both sides of the equation 14x + 2 - 14x = 12x - 14x 0 + 2 = -2x 2 = -2x Divide both sides by -2 because remember -2/-2 = 1 so it will isolate the x 2/-2 = x 2/(2(-1)) = x 1/(-1) = x -1 = x

OpenStudy (australopithecus):

\[\frac{2}{-2} = \frac{1}{-1}=\frac{1*-1}{-1*-1} = \frac{-1}{1} = -1\]

OpenStudy (australopithecus):

14x + 2 = 12x Subtract 14x from both sides of the equation 14x + 2 - 14x = 12x - 14x 0 + 2 = -2x 2 = -2x \[\frac{2}{-2} = \frac{-2x}{-2}\] \[\frac{2}{-2} = x(1)\] \[\frac{2*(-1)}{-2*(-1)} = x\] \[\frac{(1)*(-1)}{(-1)*(-1)} = x\] \[\frac{-1}{1} = x\] \[-1 = x\]